# MRGen: Diffusion-based Controllable Data Engine for MRI Segmentation towards Unannotated Modalities

### v1

@@ 4.2 Architecture
For our generation paradigm, we expect the model to leverage the large number of image-text data pairs, and the limited data with segmentation masks, to achieve controllable generation for segmentation-scarce modalities. Our proposed model (MRGen) consists of three components: (i) latent encoding; (ii) text-guided generation; and (iii) mask-conditioned generation, as detailed below.
Latent encoding. Given the high resolution of medical images, we adopt the idea of latent diffusion [42], for efficient generation within a compressed, low-dimensional latent space. As shown in Figure 2 (a), we utilize an autoencoder architecture, where a 2D MRI slice ($\mathcal{I}\in\mathbb{R}^{H\times W\times 1}$) is encoded into a latent code ($\mathbf{z}\in\mathbb{R}^{h\times w\times d}$) for diffusion training. The decoder then reconstructs the image ($\hat{\mathcal{I}}$) from the latent code, as expressed by the following:
@@
Text-guided generation. This part follows the diffusion model paradigm, consisting of a forward diffusion process and a denoising process. Concretely, the forward process progressively adds noise to the latent features ($\mathbf{z}_{0}$) over $T$ steps towards white Gaussian noise $\mathbf{z}_{T}\sim\mathcal{N}(0,1)$. At any intermediate timestep $t\in[1,T]$, the noisy features ($\mathbf{z}_{t}$) can be expressed as: $\mathbf{z}_{t}=\sqrt{\bar{\alpha}_{t}}\mathbf{z}_{0}+\sqrt{1-\bar{\alpha}_{t}}\boldsymbol{\epsilon}$, where $\boldsymbol{\epsilon}\sim\mathcal{N}(0,1)$ and $\bar{\alpha}_{t}$ represent predefined hyperparameters, which will be detailed in Section A of the Appendix.
As depicted in Figure 2 (b), the learnable denoising process typically reconstructs images from noise by estimating the noise term $\hat{\boldsymbol{\epsilon}}$. To model this process in the compressed latent space, we adopt a UNet [43]. Specifically, to generate images guided by text prompts, we design templated text prompt ($\mathcal{T}$), that consists of diverse modality labels, modality attributes, regions, and organs, for example: {mdframed}[backgroundcolor=gray!10, linewidth=0pt] “T1 MRI; fat high signal, muscle intermediate signal, water low signal, fat bright, water dark; upper abdomen; liver, spleen, and kidney”. We employ an off-the-shelf text encoder ($\phi_{\mathrm{text}}$) to encode the text prompt into embeddings, expressed as $\mathcal{C}_{T}$ = $\phi_{\mathrm{text}}(\mathcal{T})$. These text embeddings are then fed into our model via cross-attention, serving as the key and value, with visual latent codes as the query.
Mask-conditioned generation. Next, we incorporate mask conditions to enable more controllable generation. As shown in Figure 2 (c), we adopt the idea similar to ControlNet [51], i.e., initializing the mask encoder ($\phi_{\mathrm{mask}}$) with the weights of the diffusion UNet encoder, coupled with a learnable downsampling module ($\phi_{\mathrm{down}}$) to align the dimensions of the mask with the latent code. In the input mask ($\mathcal{M}\in\mathbb{R}^{H\times W\times 1}$), we use distinct intensity values to indicate the different organs, and encode it as a residual into the corresponding layers of the UNet decoder, guiding the generation process. In addition, zero convolution initialization is employed to stabilize the training process, as observed in [51]. For each layer $\mathcal{F}$ of the diffusion UNet decoder, the output $\mathbf{O}$ can now be formulated as follows:
$\mathbf{O}=\mathcal{F}(\mathbf{z}_{t})+\phi_{\mathrm{mask}}(\mathbf{z}_{t},\phi_{\mathrm{down}}(\mathcal{M}),\phi_{\mathrm{text}}(\mathcal{T}))$
@@ 4.3 Model Training
Here, we present the training procedure for our generative model, including (i) autoencoder reconstruction, (ii) text-guided pretraining, and (iii) mask-conditioned finetuning.
Autoencoder reconstruction. The autoencoder for compression is trained solely on raw images from $\mathcal{D}_{u}$, using a combination of MSE loss and KL divergence loss as follows: $\mathcal{L}_{\mathrm{VAE}}=||\mathcal{I}-\hat{\mathcal{I}}||_{2}^{2}+\gamma\mathcal{L}_{\mathrm{KL}}$, where $\mathcal{L}_{KL}$ imposes a KL-penalty towards a standard normal on the learned latent, similar to a VAE [26] and $\gamma$ represents a predefined weight.
Text-guided pretraining. The diffusion-based generative model, parameterized by $\Theta$, is trained on a large number of image-text pairs in $\mathcal{D}_{u}$, covering diverse modalities. The objective function can be formulated as the MSE loss between the added Gaussian noise ($\boldsymbol{\epsilon}$) and the prediction ($\hat{\boldsymbol{\epsilon}}$):
$\mathcal{L}=\mathbb{E}_{\mathcal{D}_{u},t\sim[1,T],\boldsymbol{\epsilon}\sim\mathcal{N}(0,1)}\Big{[}||\boldsymbol{\epsilon}-\hat{\boldsymbol{\epsilon}}(\mathbf{z}_{t},t,\mathcal{T})||_{2}^{2}\Big{]}$
This pretraining phase enables MRGen to generate radiology images across diverse modalities based on text prompts.
Mask-conditioned finetuning. The mask condition controller, consisting of the mask encoder ($\phi_{\mathrm{mask}}$) and the learnable downsampling module ($\phi_{\mathrm{down}}$), can be trained jointly on the source-domain triplet dataset, $\mathcal{D}_{s}=\{(\mathcal{I}_{s},\mathcal{T}_{s},\mathcal{M}_{s})\}$, and target-domain dataset comprising only image-text pairs, $\mathcal{D}_{t}=\{(\mathcal{I}_{t},\mathcal{T}_{t},\varnothing)\}$, with all other parameters frozen. The training objective $\mathcal{L}_{c}$ can be expressed as:
$\mathcal{L}_{c}=\mathbb{E}_{\mathcal{D}_{s},\mathcal{D}_{t},t\sim[1,T],\boldsymbol{\epsilon}\sim\mathcal{N}(0,1)}\Big{[}||\boldsymbol{\epsilon}-\hat{\boldsymbol{\epsilon}}_{c}(\mathbf{z}_{t},t,\mathcal{T},\mathcal{M})||_{2}^{2}\Big{]}$
Here, incorporating target-domain data without mask annotations prevents overfitting to the source domain.
Discussion. (i) The large-scale image-text dataset ($\mathcal{D}_{u}$) comprises a wide range of modalities, covering those of source-domain dataset ($\mathcal{D}_{s}$) and target-domain dataset ($\mathcal{D}_{t}$); (ii) Mask-conditioned finetuning on mask-annotated data allows MRGen to learn correlations between images and organ masks; (iii) Such a two-stage training strategy then enables MRGen to extend its controllable generation ability towards target domains lacking mask annotations.
Figure 3: Synthetic Data Construction Pipeline. MRGen takes text prompt and mask as conditions for controllably generating radiology images and employs a pretrained SAM2 model for automatic filtering to guarantee the quality of generated samples.
@@ 4.4 Synthetic Data for Segmentation Training
With our data engine, we aim to generate target-domain data conditioned on the segmentation masks, to facilitate training segmentation models towards unannotated modalities.
Data synethesis. At inference time, we feed the text prompt ($\mathcal{T^{\prime}}_{t}$) and organ mask ($\mathcal{M^{\prime}}_{t}$) as conditions into our MRGen model ($\Phi_{\mathrm{MRGen}}$) to generate the corresponding target-domain image ($\mathcal{I^{\prime}}_{t}$). For simplicity, when constructing the input conditions, we directly utilize region, organ information, and mask annotations from source-domain data.
Data autofilter. To ensure the fidelity of generated images to the mask conditions, we design an automatic filtering pipeline using the off-the-shelf SAM2 [41] model, as depicted in Figure 3. Specifically, we feed the mask ($\mathcal{M^{\prime}}_{t}$) and the generated image ($\mathcal{I^{\prime}}_{t}$) into SAM2 to predict a segmentation confidence score ($s_{\mathrm{conf}}$), and a pseudo mask annotation used to calculate the IoU score ($s_{\mathrm{IoU}}$) with $\mathcal{M^{\prime}}_{t}$. A sample is considered high-quality and aligned with the mask condition if both its IoU score and confidence score exceed the predefined thresholds; otherwise, it is discarded.
Source Datset Source Modality Target Dataset Target Modality Source Domain DualNorm [57] CycleGAN [58] MRGen (Ours)
CM. T1 CM. T2-SPIR 156.98 228.16 157.77 44.82
CM. T2-SPIR CM. T1 156.98 261.97 188.91 60.79
MP. T2 MP. ADC 305.56 422.73 112.82 99.38
MP. ADC MP. T2 305.56 416.31 190.82 123.55
PS. T1 PS. T2 76.95 120.36 237.52 34.35
PS. T2 PS. T1 76.95 126.27 89.26 58.90
LQ. T1 CM. T2-SPIR 109.46 281.56 182.62 88.76
CM. T2-SPIR LQ. T1 109.46 246.73 260.06 106.45
MP. ADC PR. T2 387.29 434.54 221.27 116.35
PR. T2 MP. ADC 387.29 365.10 140.72 88.43
AM. CT CM. T1 148.04 226.23 248.06 42.53
AM. CT CM. T2-SPIR 164.73 259.96 268.49 65.30
ML. CT CM. T2-SPIR 176.51 265.64 307.79 96.72
Average FID $\downarrow$ 197.06 281.20 200.47 78.95
Table 2: Quantitative Results (FID) on Generation. Smaller values indicate better performance. Here, CM., MP., PS., LQ., PR., AM., and ML., denote CHAOS-MRI, MSD-Prostate, PanSeg, LiQA, PROMISE12, AMOS22, and MSD-Liver, respectively.
Source Dataset Source Modality Target Dataset Target Modality DualNorm [57] nnUNet [21] UMamba [34]
$\mathcal{D}_{s}$ CycleGAN MRGen $\mathcal{D}_{s}$ CycleGAN MRGen Oracle Box SAM2 nnUNet [21] $\mathcal{D}_{t}$
Same Datasets, Different Modalities
CHAOS-MRI T1 CHAOS-MRI T2-SPIR 14.00 6.90 7.58 66.18 4.02 10.58 67.35 45.45 53.12 83.90
CHAOS-MRI T2-SPIR CHAOS-MRI T1 12.50 0.80 1.38 58.10 0.62 0.22 57.24 43.48 51.94 90.60
MSD-Prostate T2 MSD-Prostate ADC 1.43 5.52 40.92 57.83 0.19 45.06 52.58 61.50 65.39 82.35
MSD-Prostate ADC MSD-Prostate T2 12.94 22.20 57.06 61.95 11.80 62.00 64.05 61.07 66.40 89.80
PanSeg T1 PanSeg T2 0.21 0.68 2.40 9.78 0.38 2.13 9.34 31.69 41.11 82.28
PanSeg T2 PanSeg T1 0.11 0.30 3.59 12.07 0.27 5.08 10.29 30.94 41.18 80.27
Different Datasets, Different Modalities
LiQA T1 CHAOS-MRI T2-SPIR 19.23 16.20 7.84 31.73 11.05 8.65 37.30 62.72 70.86 70.83
CHAOS-MRI T2-SPIR LiQA T1 31.09 15.80 41.02 57.28 10.33 41.22 52.54 56.11 67.99 93.60
MSD-Prostate ADC PROMISE12 T2 1.43 17.19 42.13 35.33 23.71 43.24 37.12 50.41 53.61 87.00
PROMISE12 T2 MSD-Prostate ADC 9.84 19.20 59.95 56.88 21.75 57.21 49.77 61.50 65.39 82.35
Different Datasets, Different Modalities, CT - MRI
AMOS22 CT CHAOS-MRI T1 19.78 0.11 6.75 22.50 0.05 8.06 26.73 43.48 51.94 90.60
AMOS22 CT CHAOS-MRI T2-SPIR 16.09 8.88 52.49 56.23 3.19 43.21 60.53 45.45 53.12 83.90
MSD-Liver CT CHAOS-MRI T2-SPIR 1.58 3.12 10.14 38.67 1.65 11.06 40.93 62.72 70.86 70.83
Average DSC score 10.79 8.99 25.63 43.43 6.85 25.89 43.52 50.50 57.92 83.72
Table 3: Quantitative Results (DSC score) on Segmentation. Oracle Box denotes the DSC score compared between the perturbed oracle box prompts for SAM2, and the ground truth; $\mathcal{D}_{s}$ and $\mathcal{D}_{t}$ denote training with real and manually annotated source-domain and target-domain data respectively. Results are organized by three different cross-modality simulations, where the best results under nnUNet and UMamba are bolded respectively, while the second best are underlined.
@@ 5 Experiments
Here, we start by describing the experimental settings in Section 5.1; then, in Section 5.2 and 5.3, we present a comprehensive evaluation of our method from both quantitative and qualitative perspectives; lastly, we provide results for ablation experiments in Section 5.4 to prove the effectiveness of our proposed methods.
@@ 5.1 Experimental Settings
We aim to evaluate our proposed data engine from two aspects: (i) image generation, and (ii) segmentation. Concretely, to simulate MRI segmentation towards unannotated modalities, we construct 8 cross-modal dataset pairs within our MedGen-1M. Each pair comprises a mask-annotated source-domain dataset and an unannotated target-domain dataset. The generative model is trained on each dataset pair to synthesize target-domain samples, which are then used to train segmentation models. We assess the quality of generated images, and the performance of the segmentation models on the target-domain test set.
@@
Baselines. For image generation, we compare mask-conditioned generated images from MRGen against two other approaches: CycleGAN [58] for translating source-domain images to the target domain; and DualNorm [57] for aggressive augmentation of source-domain images. For segmentation, we evaluate models trained on five data sources: (i) source-domain data, as a baseline; (ii) DualNorm augmented data; (iii) CycleGAN translated data; (iv) MRGen generated data; and (v) manually annotated target-domain data, serving as a strong oracle baseline. We adopt nnUNet [21] and UMamba [34] as segmentation frameworks for all settings, except for DualNorm, which uses a customized UNet. Moreover, we include SAM2 [41] as a reference for interactive semi-automatic segmentation, using randomly perturbed oracle boxes as prompts.
@@
Evaluation metrics. We employ different metrics for assessment: For image generation, we adopt the commonly used Fréchet Inception Distance score (FID) [16] to assess the diversity and quality of the generated images. For the downstream segmentation models, we employ the widely used Dice Similarity Coefficient (DSC) [35] to compare the predicted segmentation mask and ground truth. Considering segmentation consistency, we stack slices back into 3D volumes for DSC calculation.
Figure 4: Qualitative Results of Controllable Generation. We present images from the source domain $\mathcal{D}_{s}$ and the target domain $\mathcal{D}_{t}$ for reference. Here, prostate, liver, right kidney, left kidney and spleen are contoured with different colors.
@@
Implementation details. All the images are resized to resolution 512 $\times$ 512 pixels. For training, we start by training the VAE with a learning rate of $5\times 10^{-5}$ and a batch size of $256$ for 50K iterations. Next, the text-guided generative model and mask condition controller are trained with a learning rate of $1\times 10^{-5}$, using batch sizes of 256 and 128 for 200K and 40K iterations, respectively. Moreover, we randomly drop text prompts with a 10% probability to enable classifier-free guidance for better performance. All training is conducted on 8×8$\times$8 × Nvidia A100 GPUs using the AdamW [32] optimizer. The downsampling factor, latent feature dimension $d$, KL divergence loss weight $\gamma$, and diffusion timesteps $T$ are set to $8$, $16$, $1\times 10^{-4}$, and $1000$, respectively. During inference, we perform 50-step sampling using DDIM [45], with the classifier-free guidance [17] weight set to $w=7.0$. We adopt the pretrained text encoder in BiomedCLIP [52] to encode text prompts, and SAM2 [41] for automatic filtering. For each conditional mask, we generate 20 candidates and select the best two that satisfy the predefined thresholds, which are set to 0.80 and 0.90 for IoU and confidence scores, respectively.
@@
Figure 5: Qualitative Results of Segmentation towards Unannotated Modalities. We present reference images from the source-domain training set $\mathcal{D}_{s}$, images from the target-domain test set $\mathcal{D}_{t}$, and corresponding predictions and ground truth. We visualize prostate on the middle row, and liver, right kidney, left kidney and spleen on the upper and lower row with different colors.

### v2

@@ 3.3 Architecture
Our research focus expects the model to leverage abundant image-text pairs and limited mask-annotated data to achieve controllable generation for underrepresented modalities. Specifically, our model (MRGen) comprises three components: (i) latent encoding; (ii) text-guided generation; and (iii) mask-conditioned generation, as detailed below.
Latent encoding. To handle high-resolution medical images, we first map them into a low-dimensional latent space for efficient training. As shown in Figure 2 (a), we employ an autoencoder, that encodes a 2D slice ($\mathcal{I}\in\mathbb{R}^{H\times W\times 1}$) into a latent representation ($\mathbf{z}\in\mathbb{R}^{h\times w\times d}$), which can be reconstructed to image ($\hat{\mathcal{I}}$) by the decoder, expressed as:
@@
To enable the generative model to effectively learn controllable generation based on texts and masks, the training process is carried out in two stages: (i) pretraining a text-guided generative model on image-text data, covering diverse modalities; and (ii) finetuning a mask condition controller jointly on data with and without mask annotations.
Text-guided generation. This part follows the diffusion model paradigm, comprising a forward diffusion process and a denoising process. Concretely, the forward process progressively adds noise to the latent features ($\mathbf{z}_{0}$) over $T$ steps towards white Gaussian noise $\mathbf{z}_{T}\sim\mathcal{N}(0,1)$. At any intermediate timestep $t\in[1,T]$, the noisy visual features ($\mathbf{z}_{t}$) is expressed as: $\mathbf{z}_{t}=\sqrt{\bar{\alpha}_{t}}\mathbf{z}_{0}+\sqrt{1-\bar{\alpha}_{t}}\boldsymbol{\epsilon}$, where $\boldsymbol{\epsilon}\sim\mathcal{N}(0,1)$, and $\bar{\alpha}_{t}$ denotes predefined hyperparameters.
As depicted in Figure 2 (b), the denoising process adopts a UNet [50] and reconstructs images from noise by estimating the noise term $\hat{\boldsymbol{\epsilon}}$. Concretely, to generate images guided by text prompts, we design templated text prompt ($\mathcal{T}$), that consists of diverse modality labels, modality attributes, regions, and organs information, for example: {mdframed}[backgroundcolor=gray!10, linewidth=0pt] “T1 MRI; fat high signal, muscle intermediate signal, water low signal, fat bright, water dark; upper abdomen; liver, spleen, and kidney”. These templated prompts ensure sufficient clinical information to distinguish distinct modalities, regions, and organs. We employ an off-the-shelf BiomedCLIP [62] text encoder ($\phi_{\mathrm{text}}$) to encode them into embeddings, denoted as $\mathcal{C}_{T}$ = $\phi_{\mathrm{text}}(\mathcal{T})$. These embeddings are integrated into our model via cross-attention, serving as the key and value, with visual features ($\mathbf{z}_{t}$) as the query. The output ($\mathbf{O}_{\mathrm{cross}}$) of each cross-attention layer ($\mathcal{F}_{\mathrm{cross}}$) is represented as:
$\mathbf{O}_{\mathrm{cross}}=\mathcal{F}_{\mathrm{cross}}(\mathbf{z}_{t},\phi_{\mathrm{text}}(\mathcal{T}))$
Mask-conditioned generation. We then incorporate mask conditions to enable more controllable generation. As presented in Figure 2 (c), we initialize the mask encoder ($\phi_{\mathrm{mask}}$) using weights from the encoder of the diffusion UNet pre-trained in the previous stage, coupled with a learnable downsampling module ($\phi_{\mathrm{down}}$) to align dimensions. The input mask ($\mathcal{M}\in\mathbb{R}^{H\times W\times 1}$) uses distinct intensity values to represent different organs, and is integrated as a residual into the UNet decoder. For each block ($\phi^{i}_{\mathrm{mask}}$) of the mask encoder, the output ($\mathbf{O}^{i}$) of the corresponding block ($\mathcal{F}^{i}$) in the diffusion UNet decoder, is formulated as:
$\mathbf{O}^{i}=\mathcal{F}^{i}(\mathbf{z}_{t})+\phi_{\mathrm{mask}}^{i}(\mathbf{z}_{t},\phi_{\mathrm{down}}(\mathcal{M}),\phi_{\mathrm{text}}(\mathcal{T}))$
@@ 3.4 Model Training
Here, we present the training procedure for our proposed model, including: (i) autoencoder reconstruction, (ii) text-guided pretraining, and (iii) mask-conditioned finetuning.
Autoencoder reconstruction. The autoencoder for compression is trained on raw images from MRGen-DB, using a combination of MSE loss and KL divergence loss as follows: $\mathcal{L}_{\mathrm{VAE}}=||\mathcal{I}-\hat{\mathcal{I}}||_{2}^{2}+\gamma\mathcal{L}_{\mathrm{KL}}$, where $\mathcal{L}_{KL}$ imposes a KL-penalty towards a standard normal on the learned latent, similar to VAE [29] and $\gamma$ denotes a predefined weight.
Text-guided pretraining. The diffusion-based generative model, parameterized by $\Theta$, is trained on a large number of image-text pairs, covering diverse modalities. The objective function is formulated as the MSE loss between the added Gaussian noise ($\boldsymbol{\epsilon}$) and the prediction ($\hat{\boldsymbol{\epsilon}}$):
$\mathcal{L}=\mathbb{E}_{t\sim[1,T],\boldsymbol{\epsilon}\sim\mathcal{N}(0,1)}\Big{[}||\boldsymbol{\epsilon}-\hat{\boldsymbol{\epsilon}}(\mathbf{z}_{t},t,\mathcal{T})||_{2}^{2}\Big{]}$
This pretraining phase enables MRGen to generate MR images across various modalities based on text prompts.
Mask-conditioned finetuning. The mask condition controller, comprising a mask encoder ($\phi_{\mathrm{mask}}$) and a downsampling module ($\phi_{\mathrm{down}}$), is jointly trained on image-text pairs with and without mask annotations, while all other parameters (including the autoencoder, text encoder, and diffusion UNet) remain frozen. The training objective $\mathcal{L}_{c}$ is:
$\mathcal{L}_{c}=\mathbb{E}_{t\sim[1,T],\boldsymbol{\epsilon}\sim\mathcal{N}(0,1)}\Big{[}||\boldsymbol{\epsilon}-\hat{\boldsymbol{\epsilon}}_{c}(\mathbf{z}_{t},t,\mathcal{T},\mathcal{M})||_{2}^{2}\Big{]}$
Here, incorporating data without mask annotations prevents the model from overfitting to those with masks.
Discussion. Such a two-stage training strategy empowers MRGen to achieve controllable generation across diverse modalities, even for those lacking mask annotations, driven by two key factors: (i) text-guided pretraining on large-scale image-text data of various modalities equips MRGen with the foundational knowledge to synthesize diverse MR images based on text prompts; (ii) mask-conditioned finetuning on partial annotated data instructs MRGen to integrate text and mask conditions, enabling controllability that generalizes to modalities included during pretraining.
Figure 3: Synthetic Data Construction Pipeline. MRGen takes text prompt and mask as conditions for controllably generating MR images and employs a pretrained SAM2 model for automatic filtering to guarantee the quality of generated samples.
@@ 3.5 Synthetic Data for Segmentation Training
With our data engine, we can then produce MR samples for training downstream segmentation models. At inference time, the text prompt ($\mathcal{T^{\prime}}$) and controlling organ mask ($\mathcal{M^{\prime}}$) are fed into our MRGen model ($\Phi_{\mathrm{MRGen}}$) as conditions to generate the corresponding MR sample ($\mathcal{I^{\prime}}$). To ensure the fidelity of the generated images on mask conditions, we design an automatic filtering pipeline using the off-the-shelf SAM2-Large [48] model, as depicted in Figure 3. Specifically, we feed the conditional mask ($\mathcal{M^{\prime}}$) and the generated image ($\mathcal{I^{\prime}}$) into SAM2 to predict a segmentation map with a confidence score ($s_{\mathrm{conf}}$), which is used to calculate the IoU score ($s_{\mathrm{IoU}}$) against $\mathcal{M^{\prime}}$. A sample is considered to be high-quality and aligned with the mask condition if both its IoU score ($s_{\mathrm{IoU}}$) and confidence score ($s_{\mathrm{conf}}$) exceed the predefined thresholds; otherwise, it is discarded.
@@ 4 Experiments
Here, we first outline the experimental settings in Sec. 4.1, followed by a comprehensive evaluation from both quantitative and qualitative perspectives in Sec. 4.2 and Sec. 4.3. Lastly, we present ablation study results in Sec. 4.4 to prove the effectiveness of our proposed method.
@@ 4.1 Experimental Settings
Unlike existing work that focuses on data augmentation for well-annotated modalities, we explore the more challenging scenarios where certain modalities lack annotations entirely, and aim to leverage generative models to synthesize data for training segmentation models towards these underrepresented modalities. Specifically, to simulate such a clinical scenario, we construct 5 cross-modality dataset pairs within our MRGen-DB, each comprising a mask-annotated source-domain dataset ($\mathcal{D}_{s}$) and an unannotated target-domain dataset ($\mathcal{D}_{t}$). Models trained on each dataset pair synthesize target-domain samples for training segmentation models. We assess our data engine from two aspects: (i) image generation quality and (ii) downstream segmentation performance on the target-domain test set.
Source Datset Source Modality Target Dataset Target Modality Source Domain DualNorm [68] CycleGAN [69] UNSB [27] MRGen (Ours)
CM. T1 CM. T2-SPIR 156.98 228.16 157.77 160.12 44.82
CM. T2-SPIR CM. T1 156.98 261.97 188.91 141.15 60.79
MP. T2 MP. ADC 305.56 422.73 112.82 303.19 99.38
MP. ADC MP. T2 305.56 416.31 190.82 346.60 123.55
PS. T1 PS. T2 76.95 120.36 237.52 80.76 34.35
PS. T2 PS. T1 76.95 126.27 89.26 76.91 58.90
LQ. T1 CM. T2-SPIR 109.46 281.56 182.62 193.05 88.76
CM. T2-SPIR LQ. T1 109.46 246.73 260.06 192.80 106.45
MP. ADC PR. T2 387.29 434.54 221.27 200.55 116.35
PR. T2 MP. ADC 387.29 365.10 140.72 252.64 88.43
Average FID $\downarrow$ 207.25 290.37 178.18 194.78 82.18
Table 2: Quantitative Results (FID) on Generation. Here, CM., MP., PS., LQ., and PR., denote CHAOS-MRI, MSD-Prostate, PanSeg, LiQA, and PROMISE12, respectively.
@@
Baselines. For generation, we compare generated images from MRGen against three representative approaches: CycleGAN [69] and UNSB [27] for translating source-domain images to the target domain; and DualNorm [68] for exhaustive augmentation of source-domain images. For segmentation, we evaluate models trained on five data sources: (i) source-domain data, as a baseline; (ii) DualNorm augmented data; (iii) CycleGAN translated data; (iv) UNSB translated data; and (v) MRGen generated data. We adopt nnUNet [23] and UMamba [40] as segmentation frameworks for all settings, except for DualNorm, which employs a customized UNet following their official codes. Comparisons with additional baselines are provided in Sec. D.2.
@@
Evaluation metrics. We employ distinct metrics for the evaluation: For image generation, we employ Fréchet Inception Distance (FID) [18] score to assess the diversity and quality of generated images. For segmentation models, we employ commonly used Dice Similarity Coefficient (DSC) [41] score to compare predicted masks with ground truth. Considering segmentation consistency, we stack slices into 3D volumes, calculate the DSC for each organ individually, and average them as the final result.
@@
Implementation details. All images are resized to 512 $\times$ 512, and training is conducted on 8×8$\times$8 × Nvidia A100 GPUs using the AdamW [38] optimizer. We start by training the autoencoder with a learning rate of $5\times 10^{-5}$ and a batch size of $256$ for 50K iterations. Next, the text-guided generative model and mask condition controller are trained with a learning rate of $1\times 10^{-5}$, using batch sizes of 256 and 128 for 200K and 40K iterations, respectively. Moreover, we randomly drop text prompts with a 10% probability to enable classifier-free guidance [19]. The compression ratio, latent dimension $d$, KL loss weight $\gamma$, and diffusion timesteps $T$ are set to $8$, $16$, $1\times 10^{-4}$, and $1000$, respectively. During inference, we perform 50-step sampling using DDIM [53], with classifier-free guidance weight $w=7.0$. For each mask, we generate 20 image candidates and select the best two satisfying the predefined thresholds, which are set to 0.80 and 0.90 for IoU and confidence scores, respectively. Conditions for target-domain data synthesis are directly derived from region, organs information, and segmentation masks of source-domain data.
Source Dataset Source Modality Target Dataset Target Modality DualNorm [68] UMamba [40] nnUNet [23]
$\mathcal{D}_{s}$ CycleGAN UNSB MRGen $\mathcal{D}_{s}$ CycleGAN UNSB MRGen
CHAOS-MRI T1 CHAOS-MRI T2-SPIR 14.00 4.02 10.58 20.56 67.35 6.90 7.58 14.03 66.18
CHAOS-MRI T2-SPIR CHAOS-MRI T1 12.50 0.62 0.22 4.41 57.24 0.80 1.38 6.44 58.10
MSD-Prostate T2 MSD-Prostate ADC 1.43 0.19 45.06 45.77 52.58 5.52 40.92 52.99 57.83
MSD-Prostate ADC MSD-Prostate T2 12.94 11.80 62.00 36.47 64.05 22.20 57.06 38.39 61.95
PanSeg T1 PanSeg T2 0.21 0.38 2.13 3.08 9.34 0.68 2.40 2.38 9.78
PanSeg T2 PanSeg T1 0.11 0.27 5.08 4.46 10.29 0.30 3.59 6.68 12.07
LiQA T1 CHAOS-MRI T2-SPIR 19.23 11.05 8.65 5.14 37.30 16.20 7.84 4.79 31.73
CHAOS-MRI T2-SPIR LiQA T1 31.09 10.33 41.22 28.52 52.54 15.80 41.02 15.28 57.28
MSD-Prostate ADC PROMISE12 T2 1.43 23.71 43.24 43.64 37.12 17.19 42.13 42.40 35.33
PROMISE12 T2 MSD-Prostate ADC 9.84 21.75 57.21 49.82 49.77 19.20 59.95 55.13 56.88
Average DSC score 10.28 8.41 27.54 24.19 43.76 10.48 26.39 23.85 44.71
Table 3: Quantitative Results (DSC score) on Segmentation. Here, $\mathcal{D}_{s}$ denotes training with manually annotated source-domain data. Results with the best and second best results are bolded and underlined, respectively.
@@
Figure 4: Qualitative Results of Controllable Generation. We present images from source domains ($\mathcal{D}_{s}$) and target domains ($\mathcal{D}_{t}$) for reference. Here, liver, right kidney, left kidney, spleen, prostate, and pancreas are contoured with different colors.